# NoisyRollout: Reinforcing Visual Reasoning with Data Augmentation

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) have strengthened the reasoning capabilities of vision-language models (VLMs). However, enhancing policy exploration to better scale test-time compute remains largely underexplored. In addition, VLMs continue to struggle with imperfect visual perception, which in turn affects the subsequent reasoning process. To this end, we propose NoisyRollout, a simple yet effective data augmentation method that mixes trajectories from both clean and moderately distorted images during RL training. By injecting targeted diversity in visual perception and the resulting reasoning patterns, NoisyRollout promotes better policy exploration through vision-oriented inductive biases, ultimately leading to more robust reasoning behaviors. We further adopt a noise annealing schedule that gradually reduces distortion strength over training, leveraging noisy signals early on while ensuring training stability in later stages. Crucially, our method is easy-to-adopt—requiring no additional training cost and no modifications to the RL objective. Extensive experiments on $\mathbf{2}$ distinct training datasets demonstrate that NoisyRollout achieves state-of-the-art performance among open-source RL-tuned models across $\mathbf{5}$ out-of-domain reasoning and perception benchmarks. Furthermore, we validate the effectiveness of NoisyRollout across model sizes ($7$B and $32$B) and data scales (from $1$K to $6$K), highlighting its generalizability and scalability.
@@ 1 Introduction
However, scaling test-time compute via RL requires more than sheerly generating longer outputs [40], and VLMs face unique challenges in this process. A key challenge is effective policy exploration, enabling policies to discover behaviors that generalize well beyond training data [76, 72]—an area largely underexplored in VLM research. Traditional practices, such as increasing rollout temperature to promote decoding diversity [78], often introduce superficial variability without meaningfully directing policies toward more robust or informative behaviors. Moreover, VLMs inherently struggle with imperfect visual perception [35, 68], which negatively impacts subsequent reasoning processes [82, 86, 25]. Despite this, recent efforts [41, 48, 13] tend to adapt RL methods directly from the LLM domain. Such approaches often fail to address these perceptual challenges, thereby hindering the efficient development of visual reasoning capabilities through RL.
Figure 1: Accuracy improvement over Qwen2.5-VL-7B-Instruct on $5$ out-of-domain benchmarks, covering both visual reasoning tasks (from MathVerse to WeMath) and a visual perception task (HallusionBench). Both Qwen2.5-VL-GRPO-7B and NoisyRollout-7B are fine-tuned by ourselves (denoted with $\ddagger$) using vanilla GRPO with only $\mathbf{2.1}$K training samples from Geometry3K. The exact accuracy of NoisyRollout-7B is annotated above each corresponding bar in parentheses.
Tackling the challenges of policy exploration and perceptual limitations in VLMs during RL training, we propose NoisyRollout, a simple yet powerful data augmentation technique for VLMs that introduces meaningful rollout diversity. Specifically, for each training sample consisting of an input image $I$ and a corresponding text query $\mathbf{q}$, the old policy ($\pi_{\theta_{\mathrm{old}}}$) produces two sets of rollouts based on the original clean image and a moderately distorted version of the same image, respectively. While the current policy ($\pi_{\theta}$) is updated solely by conditioning on the clean image and text query pair $(I,\mathbf{q})$, the two sets of rollouts form a group, collectively contributing to computing the reward baseline and normalized advantage in Group Relative Policy Optimization (GRPO) [56]. This hybrid rollout strategy enables the policy to achieve more targeted and efficient exploration, ultimately leading to more robust visual reasoning via RL through two key mechanisms:
@@
We conduct extensive experiments to validate the effectiveness of NoisyRollout. Trained with only $2.1$K samples from the Geometry3K [42] dataset using Qwen2.5-7B-VL-Instruct [4], Figure 1 shows that NoisyRollout achieves superior performance across $5$ out-of-domain visual reasoning and perception benchmarks [43, 18, 82, 64, 52] (MathVerse $53.2$%, MathVision $28.5$%, and HallusionBench $72.1\%$). It outperforms both open-source RL-tuned models [48, 67] and those utilizing large-scale supervised fine-tuning (SFT) before RL [73, 81, 13]. Furthermore, it consistently surpasses its direct baseline (vanilla GRPO) on both in-domain and out-of-domain tasks, all within a fixed total rollout budget. Crucially, these out-of-domain improvements generalize across different model sizes (e.g., $7$B to $32$B) as well as training corpora and data scales (e.g., MMK12 [48] with $1$K to $6$K samples). These empirical results, combined with its simplicity and lightweight characteristics, establish NoisyRollout as a potentially scalable approach. We summarize the contributions of this paper as follows:
  • We identify and investigate critical, yet underexplored, challenges in policy exploration and perceptual robustness that arise during the RL training of VLMs.
  • We introduce NoisyRollout, advancing a data augmentation perspective to improve visual reasoning generalization, by mixing trajectories from both clean and moderately distorted images.
  • Extensive experiments demonstrate NoisyRollout’s consistent outperformance of vanilla GRPO across model sizes and data scales, achieving state-of-the-art results among open-source RL-tuned VLMs—without additional training overhead or complex system modifications.
Figure 2: Illustration of the NoisyRollout workflow. Solid lines depict the generation and use of clean rollouts from the clean (original) input $(I,\mathbf{q})$, while dashed lines depict the generation and use of noisy rollouts from the corresponding noisy input $(\tilde{I},\mathbf{q})$. The distorted image $\tilde{I}$ is obtained by applying a distortion function $\tilde{I}=T_{\alpha_{t}}(I)$ with distortion strength $\alpha_{t}$. The distortion level $\alpha_{t}$ is controlled by a noise annealing schedule, which gradually decreases distortion during training. The terms $\{\mathbf{o}_{i}\}_{i=1}^{n_{1}+n_{2}}$, $\{r_{i}\}_{i=1}^{n_{1}+n_{2}}$, and $\{A_{i}\}_{i=1}^{n_{1}+n_{2}}$ represent mixed trajectories, rewards, and advantages, respectively. Notably, policy optimization conditions only on clean inputs, while the corresponding noisy inputs are used solely to collect noisy rollouts.
@@ 2 NoisyRollout: A Free-Lunch with Noisy Reinforcement Learning
[…]
@@
3 Experiments
Dataset. We use EasyR1 [85] as our reinforcement learning training framework, which is built on verl [57] and specifically designed for VLMs. Our experiments utilize two datasets: Geometry3K [42], focused on geometric problem solving, and MMK12 [48], covering diverse K-12 math topics. These datasets comprise $2.1$K and $6.4$K training samples respectively. We processed them by converting all questions from multiple-choice to free-form format to prevent reward hacking and model guessing.
@@ 3 Experiments
InternVL-2.5-8B-Instruct [9] - 39.5 19.7 64.4 - 67.3$\dagger$
LLaVA-OneVision-7B [31] - 26.2 - 63.2 - 48.4$\dagger$
Kimi-VL-16B [27] - 44.9 21.4 68.7 - 66.2$\dagger$
@@
Implementation details. Following prior work [48, 67], we initialize our policy models with Qwen2.5-VL-7/32B-Instruct, which exhibit strong foundational capabilities well-suited for subsequent RL training. All experiments are conducted using $8$ A100 GPUs (40G for $7$B model, 80G for $32$B model). We keep the vision encoder frozen for training stability and parameter efficiency. For other general RL-related hyperparameters, we adopt the default settings from EasyR1: a global batch size of $128$, a rollout batch size of $512$, a rollout temperature of $1.0$, and a learning rate of $1\text{e}{-}6$. To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy. [^3] The code implementation can be found at verl. For NoisyRollout-specific configurations, we adopt Gaussian noise as the image distortion strategy, and apply a sigmoid-shaped annealing schedule:
@@
where $\gamma$ determines the midpoint of the annealing curve and $\lambda$ controls its steepness. Figure 7 illustrates the visual effects of applying different levels of Gaussian noise to a clean image. We defer the discussion of image distortion strategies (e.g., cropping, rotation), noise annealing strategies (e.g., power, exponential), and proportions of noisy rollouts in total rollouts to Appendix A. The reasoning and direct-answer templates used in our experiments are shown in Appendix I. Additional implementation details regarding the number of training steps/epochs and the hyperparameters for image distortion and noise annealing are presented in Appendix K.
@@ 3.1 Main Results
Moreover, while Qwen2.5-7B-VL-Instruct’s perception accuracy on HallusionBench drops from $71.2$% to $64.6$% when switching from direct-answer to reasoning templates, [^4] This degradation caused by the reasoning template has also been observed in previous studies [6, 25]. NoisyRollout achieves $72.1$% with the reasoning prompt (compared to vanilla GRPO’s $69.8$%). The final subplot in Figure 3 further confirms that NoisyRollout enhances perception quality during reasoning, achieving a higher Bradley–Terry win rate over vanilla GRPO (See Appendix D for details). These results indicate that our hybrid rollout strategy enhances visual perception by promoting better policy exploration through vision-oriented inductive biases.
Result 2: Sample efficiency. NoisyRollout demonstrates exceptional data efficiency by generalizing with only 2.1K training samples from Geometry3K, whereas comparable models require significantly more data or even additional SFT as warm-up training. For example, Table 1 indicates that OpenVLThinker-7B needs 35K SFT samples and 15K RL samples but reaches only $48.0\%$ on MathVerse and $71.5$% on MathVista. This efficiency stems from NoisyRollout’s use of noisy training signals that foster targeted exploration during RL, enabling effective generalization from limited samples.
Result 3: Robustness across training datasets and model sizes. NoisyRollout consistently improves upon vanilla GRPO, demonstrating strong robustness across model sizes and training datasets. As shown in Tables 1 and 2, the $7$B model trained on MMK12 achieves gains of $1.2$%, $1.3$%, and $1.9$% over GRPO on MathVerse, MathVista, and HallusionBench, respectively. Similarly, the 32B model trained on MMK12 surpasses GRPO by $0.4$%, $0.7$%, and $1.1$% on the same benchmarks.
Notably, in certain benchmarks like MathVerse and MathVista, the performance gains of NoisyRollout over vanilla GRPO are smaller for the $32$B model than for the $7$B model. This is likely because the $32$B model’s initial policy was already fine-tuned via RL, [^5] https://qwenlm.github.io/blog/qwen2.5-vl-32b/ whereas the 7B model’s was not.
Table 3: Performance comparison under different rollout temperature settings, with the total number of rollouts fixed at $12$. In vanilla GRPO, “$n(6):1.0$, $n(6):1.2$” indicates $6$ rollouts with temperature $1.0$ and another $6$ with temperature $1.2$. In NoisyRollout, “$n_{1}(6):1.0$” denotes $6$ rollouts per sample generated from clean input $(I,\mathbf{q})$ with temperature $1.0$, while “$n_{2}(6):1.0$” denotes $6$ rollouts per sample from noisy input $(\tilde{I},\mathbf{q})$ with temperature $1.0$. “Geo3K” represents the test set of Geometry3K dataset. “Avg.” represents average accuracy (%) across six benchmarks.
@@
Figure 4: Comparison of accuracy and diversity metrics (%) across RL training steps (00 to $40$). The left two subfigures contrast NoisyRollout versus vanilla GRPO (both with temperature $1.0$), while the right two demonstrate the effects of different temperature settings ($0.8$, $1.0$, $1.2$) on vanilla GRPO.
@@ 3.2 Ablation Study: More Effective Rollout Diversity with Noisy Trajectories
3.3 Ablation Study: Impact of Hyperparameters and Module Design
Data scale. Although Geometry3K is a high-quality training dataset, its limited size ($2.1$K samples) prevents a thorough investigation of the scaling behavior of NoisyRollout compared to vanilla GRPO. To enable such analysis, we additionally consider MMK12, which contains $6.4$K samples after preprocessing. Figure 6 shows NoisyRollout consistently outperforms vanilla GRPO across various data scales, ranging from $1.1$K to $6.4$K. Notably, the performance gains do not diminish as the dataset size increases, suggesting that NoisyRollout has strong potential for use in large-scale training regimes.
Table 5: Ablation study on the noise annealing strategy.
Method Geometry3K OOD Avg.
Qwen2.5-VL-7B-Instruct 39.4 53.3
+ Vanilla GRPO 51.4 57.2
+ NoisyRollout w.o. Noise Annealing 43.9 58.0
+ NoisyRollout 54.9 59.2
@@ 3.3 Ablation Study: Impact of Hyperparameters and Module Design
Initial noise step. We evaluate the impact of noise strength by varying the initial Gaussian noise step, as shown in Table 4. Gradually increasing the initial noise step $\alpha_{0}$ from 00 to $500$ consistently improves performance across all evaluation categories, suggesting that moderate noise promotes exploration and enriches the training signal. However, exceeding this threshold leads to performance degradation, as overly distorted images (see Figure 7) yield noisy rollouts with average near-zero rewards. These excessively noisy samples introduce harmful distribution shifts during policy updates, ultimately destabilizing the learning process. Additional ablation results on the MMK12 dataset are deferred to Appendix A. [^7] We also include additional ablations (e.g., number of rollouts and data seed variations) in Appendix A.

### v4

@@ Abstract
Recent advances in reinforcement learning (RL) have strengthened the reasoning capabilities of vision-language models (VLMs). However, enhancing policy exploration to better scale test-time compute remains largely underexplored. In addition, VLMs continue to struggle with imperfect visual perception, which in turn affects the subsequent reasoning process. We introduce NoisyRollout, a simple yet effective data augmentation method that addresses these issues by mixing training trajectories from both clean and moderately distorted images. This approach injects perceptual diversity, encouraging better policy exploration and leading to more robust reasoning. A noise annealing schedule gradually reduces distortion strength, aiding exploration early in training while ensuring later stability. Crucially, our method is easy-to-adopt—requiring no additional training cost and no modifications to the RL objective. Extensive experiments on $\mathbf{2}$ distinct training datasets demonstrate that NoisyRollout achieves state-of-the-art performance among open-source RL-tuned models across $\mathbf{5}$ out-of-domain reasoning and perception benchmarks. Furthermore, we validate the effectiveness of NoisyRollout across model sizes (77B and 3232B), data scales (from 11K to 66K) and image augmentation types (Gaussion noise and rotation), highlighting its generalizability and scalability. Refer to caption Figure 1: An Illustration of the NoisyRollout workflow. Solid lines depict the generation and use of clean rollouts from the clean (original) input $(I,\mathbf{q})$, while dashed lines depict the generation and use of noisy rollouts from the corresponding noisy input $(\tilde{I},\mathbf{q})$. The distorted image $\tilde{I}$ is obtained by applying a distortion function $\tilde{I}=T_{\alpha_{t}}(I)$ with distortion strength $\alpha_{t}$. The distortion level $\alpha_{t}$ is controlled by a noise annealing schedule, which gradually decreases distortion during training. Rollouts from both sources are mixed to form the final trajectories $\{\mathbf{o}_{i}\}_{i=1}^{n_{1}+n_{2}}$, rewards $\{r_{i}\}_{i=1}^{n_{1}+n_{2}}$, and advantages $\{A_{i}\}_{i=1}^{n_{1}+n_{2}}$. Crucially, policy optimization conditions only on clean inputs $(I,\mathbf{q})$; the corresponding noisy inputs $(\tilde{I},\mathbf{q})$ are used solely to collect diverse rollouts for exploration.
@@ 1 Introduction
However, scaling test-time compute via RL requires more than sheerly generating longer outputs [45], and VLMs face unique challenges in this process. A key challenge is effective policy exploration, enabling policies to discover behaviors that generalize well beyond training data [87, 82]—an area largely underexplored in VLM research. Traditional practices, such as increasing rollout temperature to promote decoding diversity [89], often introduce superficial variability without meaningfully directing policies toward more robust or informative behaviors.
Moreover, VLMs inherently struggle with imperfect visual perception [40, 78], which negatively impacts subsequent reasoning processes [93, 97, 27]. Despite this, recent efforts [46, 53, 13] tend to adapt RL methods directly from the LLM domain. Such approaches often fail to take these perceptual challenges into consideration, thereby hindering the efficient development of visual reasoning capabilities through RL training.
Figure 2: Accuracy improvement over Qwen2.5-VL-7B-Instruct on 55 out-of-domain benchmarks, covering both visual reasoning tasks (from MathVerse to WeMath) and a visual perception task (HallusionBench). Both Qwen2.5-VL-GRPO-7B and NoisyRollout-7B are fine-tuned by ourselves (denoted with $\ddagger$) using vanilla GRPO with only $\mathbf{2.1}$K training samples from Geometry3K. The exact accuracy of NoisyRollout-7B is annotated above each corresponding bar in parentheses.
Tackling the challenges of policy exploration and perceptual limitations in VLMs during RL training, we propose NoisyRollout, a simple yet powerful data augmentation technique for VLMs that introduces meaningful rollout diversity. Specifically, for each training sample consisting of an input image $I$ and a corresponding text query $\mathbf{q}$, the old policy ($\pi_{\theta_{\mathrm{old}}}$) produces two sets of rollouts based on the original clean image and a moderately distorted version of the same image, respectively.
While the current policy ($\pi_{\theta}$) is updated solely by conditioning on the clean image and text query pair $(I,\mathbf{q})$, the two sets of rollouts form a group, collectively contributing to computing the reward baseline and normalized advantage in Group Relative Policy Optimization (GRPO) [64]. This hybrid rollout strategy enables the policy to achieve more targeted and efficient exploration, ultimately leading to more robust visual reasoning via RL through two key mechanisms:
@@
We conduct extensive experiments to validate the effectiveness of NoisyRollout. Trained with only 2.12.1K samples from the Geometry3K [47] dataset using Qwen2.5-7B-VL-Instruct [4], Figure 2 shows that NoisyRollout achieves superior performance across $5$ out-of-domain visual reasoning and perception benchmarks [48, 19, 93, 74, 57] (MathVerse 53.253.2%, MathVision 28.528.5%, and HallusionBench $72.1\%$). It outperforms both open-source RL-tuned models [53, 77] and those utilizing large-scale supervised fine-tuning (SFT) before RL [83, 92, 13]. Furthermore, it consistently surpasses its direct baseline (vanilla GRPO) on both in-domain and out-of-domain tasks, all within a fixed total rollout budget. Crucially, these out-of-domain improvements generalize across different model sizes (e.g., 77B to 3232B) as well as training corpora and data scales (e.g., MMK12 [53] with 11K to 66K samples). These empirical results, combined with its simplicity and lightweight characteristics, establish NoisyRollout as a potentially scalable approach.
@@ 2 NoisyRollout: A Free-Lunch with Noisy Reinforcement Learning
$\Biggl[\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}\min\ \!\Biggl(\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})}\hat{A}_{i},\mathrm{clip}\ \!\Bigl(\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})},\,1-\epsilon,\,1+\epsilon\Bigr)\hat{A}_{i}\Biggr)\Biggr].$ (2)
Algorithm 1 NoisyRollout: Noisy Reinforcement Fine-Tuning
2:for $t=1$ to $t_{\text{max}}$ do
3:  Sample batch $(I,\mathbf{q})\sim p_{\mathcal{D}}$
4:  Set noise strength $\alpha_{t}=\eta(\alpha_{0},t,t_{\text{max}})$ $\triangleright$ Annealing schedule
5:  Generate distorted images $\tilde{I}=T_{\alpha_{t}}(I)$
6:  Sample $\{\mathbf{o}_{j}\}_{j=1}^{n_{1}}$ from $\pi_{\theta_{\text{old}}}(\mathbf{o}\mid I,\mathbf{q})$ $\triangleright$ Clean rollouts
7:  Sample $\{\mathbf{o}_{k}\}_{k=n_{1}+1}^{n_{1}+n_{2}}$ from $\pi_{\theta_{\text{old}}}(\mathbf{o}\mid\tilde{I},\mathbf{q})$ $\triangleright$ Noisy rollouts
10:  Update policy using:
12:  $\theta\leftarrow\theta-\nabla_{\theta}\mathcal{J}(\theta)$ $\triangleright$ Update conditioned on clean images only
13:  $\theta_{\text{old}}\leftarrow\theta$ $\triangleright$ Update old policy parameters
14:end for
@@
[…]
@@
3 Experiments
Dataset. We use EasyR1 [96] as our reinforcement learning training framework, which is built on verl [65] and specifically designed for VLMs. Our experiments utilize two datasets: Geometry3K [47], focused on geometric problem solving, and MMK12 [53], covering diverse K-12 math topics. These datasets comprise 2.12.1K and 6.46.4K training samples respectively. We processed them by converting all questions from multiple-choice to free-form format to prevent reward hacking and model guessing.
Implementation details. Following prior work [53, 77], we initialize our policy models with Qwen2.5-VL-7/32B-Instruct, which exhibit strong foundational capabilities well-suited for subsequent RL training. All experiments are conducted using $8$ A100 GPUs (40G for 77B model, 80G for 3232B model). We keep the vision encoder frozen for training stability and parameter efficiency. For other general RL-related hyperparameters, we adopt the default settings from EasyR1: a global batch size of $128$, a rollout batch size of $512$, a rollout temperature of $1.0$, and a learning rate of $1\text{e}{-}6$. To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy. [^3] The code implementation can be found at verl. For NoisyRollout-specific configurations, we adopt Gaussian noise as the default image distortion strategy, and apply a sigmoid-shaped annealing schedule:
@@ 3 Experiments
where $\gamma$ determines the midpoint of the annealing curve and $\lambda$ controls its steepness. Figure 7 illustrates the visual effects of applying different levels of Gaussian noise to a clean image. We defer the discussion of unsuccessful image distortion strategies (e.g., cropping), noise annealing strategies (e.g., power, exponential), and proportions of noisy rollouts in total rollouts to Appendix A. The reasoning and direct-answer templates used in our experiments are shown in Appendix H. Additional implementation details regarding the number of training steps/epochs and the hyperparameters for image distortion and noise annealing are presented in Appendix J.
@@ 3.1 Main Results
Moreover, while Qwen2.5-7B-VL-Instruct’s perception accuracy on HallusionBench drops from 71.271.2% to 64.664.6% when switching from direct-answer to reasoning templates, [^4] This degradation caused by the reasoning template has also been observed in previous studies [6, 27]. NoisyRollout achieves 72.172.1% with the reasoning prompt (compared to vanilla GRPO’s 69.869.8%). The final subplot in Figure 3 further confirms that NoisyRollout enhances perception quality during reasoning, achieving a higher Bradley–Terry win rate over vanilla GRPO (See Appendix C for details). These results indicate that our hybrid rollout strategy enhances visual perception by promoting better policy exploration through vision-oriented inductive biases.
@@
Result 2: Sample efficiency. NoisyRollout demonstrates exceptional data efficiency by generalizing with only 2.1K training samples from Geometry3K, whereas comparable models require significantly more data or even additional SFT as warm-up training. For example, Table 1 indicates that OpenVLThinker-7B needs 35K SFT samples and 15K RL samples but reaches only $48.0\%$ on MathVerse and 71.571.5% on MathVista. This efficiency stems from NoisyRollout’s use of noisy training signals that foster targeted exploration during RL, enabling effective generalization from limited samples.
Result 3: Robustness across training datasets and model sizes. NoisyRollout consistently improves upon vanilla GRPO, demonstrating strong robustness across model sizes and training datasets. As shown in Tables 1 and 2, the 77B model trained on MMK12 achieves gains of 1.21.2%, 1.31.3%, and 1.91.9% over GRPO on MathVerse, MathVista, and HallusionBench, respectively. Similarly, the 32B model trained on MMK12 surpasses GRPO by 0.40.4%, 0.70.7%, and 1.11.1% on the same benchmarks.
Notably, in certain benchmarks like MathVerse and MathVista, the performance gains of NoisyRollout over vanilla GRPO are smaller for the 3232B model than for the 77B model. This is likely because the 3232B model’s initial policy was already fine-tuned via RL, [^5] https://qwenlm.github.io/blog/qwen2.5-vl-32b/ whereas the 7B model’s was not.
Table 3: Performance comparison under different rollout temperature settings, with the total number of rollouts fixed at $12$. In vanilla GRPO, “$n(6):1.0$, n(6):1.2n(6):1.2” indicates 66 rollouts with temperature $1.0$ and another 66 with temperature $1.2$. In NoisyRollout, $n_{1}(6):1.0$” denotes 66 rollouts per sample generated from clean input $(I,\mathbf{q})$ with temperature $1.0$, while $n_{2}(6):1.0$” denotes 66 rollouts per sample from noisy input $(\tilde{I},\mathbf{q})$ with temperature $1.0$. “Geo3K” represents the test set of Geometry3K dataset. “Avg.” represents average accuracy (%) across six benchmarks.
@@ 3.2 Ablation Study: More Effective Rollout Diversity with Noisy Trajectories
Figure 4: Comparison of accuracy and diversity metrics (%) across RL training steps (0 to $40$). The left two subfigures contrast NoisyRollout versus vanilla GRPO (both with temperature $1.0$), while the right two demonstrate the effects of different temperature settings ($0.8$, $1.0$, $1.2$) on vanilla GRPO.
3.3 Ablation Study: Impact of Hyperparameters and Module Design
@@
Table 4: Ablation study on the noise annealing strategy.
Method Geometry3K OOD Avg.
Qwen2.5-VL-7B-Instruct 39.4 53.3
+ Vanilla GRPO 51.4 57.2
+ NoisyRollout w.o. Noise Annealing 43.9 58.0
+ NoisyRollout 54.9 59.2
Table 5: Empirical validation with additional augmentation types.
Method Augmentation MathVerse MathVision MathVista WeMath HallusionBench Average
GRPO None 50.8 27.3 70.5 67.4 69.8 57.2
NoisyRollout Gaussian noise 53.2 (+2.4) 28.5 (+1.2) 72.6 (+2.1) 69.6 (+2.2) 72.1 (+2.3) 59.2 (+2.0)
NoisyRollout Rotation 52.5 (+1.7) 28.1 (+0.8) 71.9 (+1.4) 68.1 (+0.7) 70.2 (+0.4) 58.2 (+1.0)
Image augmentation type. To validate that the benefits of NoisyRollout are not limited to a single type of distortion, we evaluate its performance using rotation as a representative geometric transformation, in addition to our default Gaussian noise. The results, presented in Table 5, show that both augmentation strategies outperform the vanilla GRPO baseline. Specifically, using rotation boosts the average score from 57.257.2% to 58.258.2%. While Gaussian noise yields superior results with an average score of 59.259.2%, the meaningful gains from rotation demonstrate the generalizability of our approach. This suggests that the core mechanism of NoisyRollout —enhancing policy exploration through diverse visual inputs—is robust and effective across different augmentation techniques.
@@
Data scale. Although Geometry3K is a high-quality training dataset, its limited size (2.12.1K samples) prevents a thorough investigation of the scaling behavior of NoisyRollout compared to vanilla GRPO. To enable such analysis, we additionally consider MMK12, which contains 6.46.4K samples after preprocessing. Figure 6 shows NoisyRollout consistently outperforms vanilla GRPO across various data scales, ranging from 1.11.1K to 6.46.4K. Notably, the performance gains do not diminish as the dataset size increases, suggesting that NoisyRollout has strong potential for use in large-scale training regimes.
@@
Initial noise step. We evaluate the impact of noise strength by varying the initial Gaussian noise step, as shown in Table 6. Gradually increasing the initial noise step $\alpha_{0}$ from 0 to $500$ consistently improves performance across all evaluation categories, suggesting that moderate noise promotes exploration and enriches the training signal. However, exceeding this threshold leads to performance degradation, as overly distorted images (see Figure 7) yield noisy rollouts with average near-zero rewards. These excessively noisy samples introduce harmful distribution shifts during policy updates, ultimately destabilizing the learning process. Additional ablation results on the MMK12 dataset are deferred to Appendix A. [^7] We also include additional ablations (e.g., number of rollouts and data seed variations) in Appendix A.